OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 160r^4 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of sqrt(r^7) square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (anonymous):

i gotta say that this isn't much of a collage :P total area / small area = total pieces (160r^4) / (sqrt(r^7)) \[\frac{160r^4}{\sqrt{r^7}} = \frac{160r^4*\sqrt{r^7}}{r^7}= \frac{160\sqrt{r^7}}{r^3}\] cheers